Direction: Solve the following problems below. Show your solutions.



a. A charge of 3x10^-8 C experiences an electrostatic force of 6.0x10^-8 N. Compute the force per coulomb that the charge experiences.



b. Compute the electric field experienced by a negative test charge given the following conditions in a vacuum.


i. Source charge: -4.33x10^-12 C


ii. Distance from source charge: 2.33x10^-5 m

1
Expert's answer
2022-02-02T05:15:27-0500

(a) the force per coulomb

"E=F/q=\rm 6.0*10^{-8}\:N/3.0*10^{-8}\: C=2.0\; N/C"

(b) electric field experienced by a negative test charge

"E=kq/r^2\\
=9*10^9*4.33*10^{-12}/(2.33*10^{-5})^2\\
=7.19*10^7\:\rm N/C"